Circular No. 6711

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997da IN IC 1216 M. Schwartz reports his discovery of an apparent supernova (mag ∼

17) on a CCD image taken with the 0.35-m Tenagra Observatory patrol telescope on July 31.399 UT. SN 1997da is located 3′′ west and 6′′ south of the nucleus of IC 1216. G. V. Williams, Smithsonian Astrophysical Observatory, provides the following measurement for SN 1997da from a GIF image supplied by Schwartz: α = 16^{h}15^{m}54^{s}.96, δ = +68^{o}20′53′′.3 (equinox 2000.0; uncertainty ∼ 1′′). A. V. Filippenko, D. C. Leonard, and A. M. Gilbert, University of California at Berkeley, report that CCD spectra (range 330–1000 nm) obtained with the Lick 3-m Shane reflector on Aug. 4 show this to be a supernova of type II, probably within 1.5 months after outburst; Hα has a P-Cyg profile, and there are some Fe II lines.

SUPERNOVA 1997db IN UGC 11861 Schwartz also reports his discovery of an apparent supernova on a

CCD image taken on Aug. 2.30 UT. The following position and magnitude were measured by W. Offutt (Cloudcroft, NM) from an image supplied by Schwartz: α = 21^{h}56^{m}20^{s}.53, δ = +73^{o}14′49′′.4 (equinox 2000.0), V = 16.9. Note that SN 1995ag was in the same galaxy (IAUC 6244).

Filippenko et al. report that CCD spectra (range 330–1000 nm) obtained with the 3-m Shane reflector on Aug. 4 show that SN 1997db is a supernova of type II, many months past outburst and entering the nebular phase. There are prominent emission lines of [Ca II] 730-nm and the Ca II near-infrared triplet, as well as very strong Hα. SN 1997db is located ∼ 14′′ west and 50′′ south from the ill-defined host-galaxy nucleus.

CEPHEUS X-4 R. W. Argyle, Royal Greenwich Observatory, reports an accurate op-

tical position for the suggested counterpart to Cepheus X-4 (IAUC 6698) obtained with the Carlsberg Automatic Meridian Circle on La Palma. The position from 8 observations for the International Celestial Reference Frame is: α = 21^{h}39^{m}30^{s}.630 ± 0^{s}.012, δ = +56^{o}59′09′′.99 ± 0′′.12 (equinox J2000.0; epoch 1997.55). Photometry is unreliable, but the object certainly shows signs of variability.

1997 August 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green